OpenStudy (anonymous):

a graphic designer is designing a layout on a rectangular page that will contain 24 square inches of print. The margins at the top and bottom of the page will be 1.5 inches wide, and the margins on each side will be 1 inch wide. What page dimensions will minimize the amount of paper used?

OpenStudy (anonymous):

Call the printed height a, width b then ab is 24 and total area is (a+2)(b+3). Now just sub in using ab =24 and differentiate for min.

OpenStudy (anonymous):

U know what to do?

OpenStudy (anonymous):

i get the (a+2)(b+3) part....i don't know if i'm doing the rest correctly

OpenStudy (anonymous):

Just sub for a or b using ab =24 and u get an equation all in a or b. And diff it like the other one, set = 0 for min/max

OpenStudy (anonymous):

okay i used x and y instead of a and b and used all y: 24 = (x+2)(y+3) so x = 18-2y/(y+3) and then plug that in for x in A = xy to get -2y^2-6y-72/(y+3) .....is this correct so far?

OpenStudy (anonymous):

The total area is say A = (x+2)(y+3) with xy = 24 so say y =24/x. U don't need the 24 except as a relation between x and y.

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

u r welcome.